OpenStudy (anonymous):

Mark, Jessica and Nate each downloaded music from the same website. Mark downloaded 10 songs in total consisting of pop, rock, and hip hop. Jessica downloaded five times as many pop songs, twice as many rock and three times as many hip hop songs as Mark. She downloaded 28 songs total. If Nate downloaded 20 songs total with three times as many pop songs, three times as many rock songs, and the same number of hip hop songs as Mark, how many songs of each type did Mark download?

OpenStudy (anonymous):

A.) Mark downloaded 2 pop songs, 5 rock songs and 3 hip hop songs. B.) Mark downloaded 3 pop songs, 2 rock songs and 5 hip hop songs. C.) Mark downloaded 4 pop songs, 2 rock songs and 4 hip hop songs. D.) Mark downloaded 1 pop songs, 4 rock songs and 5 hip hop songs.

OpenStudy (anonymous):

Is there anyone that can help me with this problem?

OpenStudy (anonymous):

@LouJoeLou7 - mate i'd love to , but i gtg . sorry again .

OpenStudy (anonymous):

Ah okay.

OpenStudy (anonymous):

I think we think that the system of equations is- x+y+z=10 5x+2y+3z=28 3x+3y+z=20

OpenStudy (anonymous):

Whoops, I think that*

OpenStudy (anonymous):

.

OpenStudy (anonymous):

Thx everyone for the help... Got my answer on yahoo instead lol.